Write the expression as a single logarithm.


4log6x-1/5 log6z + 2log6y



1
Expert's answer
2022-03-13T18:15:42-0400

"\frac{4log_{6}(x-1)}{5log_{6}z+2log_{6}y}"


"=>\frac{log_{6}(x-1)⁴}{log_{6}z⁵+log_{6}y²}"


"=>\frac{log_{6}(x-1)⁴}{log_{6}z⁵y²}"


Recall that "\frac{log_{x} a}{log_{x}b} = log_{b}a"


"=>\frac{log_{6}(x-1)⁴}{log_{6}z⁵y²}={log_{z⁵y}(x-1)⁴}"